# Local stability guarantees for data-driven quadratically nonlinear models

### v1

@@ II.4 Analytical estimates of the local stability of linear-quadratic systems
The time derivative of the Lyapunov matrix given by Eq. (II.4) is
$\dot{v}(\bm{y})$ $=\bm{y}^{T}\bm{P}\dot{\bm{y}}$ (27a)
$=\bm{d}^{T}\bm{P}\bm{y}+\bm{y}^{T}\bm{P}\bm{A}^{S}\bm{y}+\bm{y}^{T}\bm{P}\bm{H}(\bm{y\otimes y})$ (27b)
$=\bm{\tilde{d}}^{T}\bm{y}+\bm{y}^{T}\bm{\tilde{A}}^{S}\bm{y}+\bm{y}^{T}\bm{\tilde{H}}(\bm{y\otimes y}).$ (27c)
Here $\bm{\tilde{d}}=\bm{P}\bm{d}$, $\bm{\tilde{A}}^{S}=\bm{P}\bm{A}^{S}$ and $\bm{\tilde{H}}=\bm{P}\bm{H}=\bm{P}\mathcal{T}(\bm{Q})$. We can then introduce $\tilde{Q}^{\prime}_{ijk}=\tilde{Q}_{ijk}+\tilde{Q}_{jki}+\tilde{Q}_{kij}$, and $\bm{H}_{0}=\mathcal{T}(\tilde{\bm{Q}}^{\prime})$ so that $\dot{v}(\bm{y})$ now is
$\dot{v}(\bm{y})$ $=\bm{\tilde{d}}^{T}\bm{y}+\bm{y}^{T}\bm{\tilde{A}}^{S}\bm{y}+\frac{1}{3}\bm{y}^{T}\bm{H}_{0}(\bm{y}\otimes\bm{y})$ (28a)
$\leq\|\bm{\tilde{d}}^{T}\bm{y}\|_{2}+\lambda_{1}\|\bm{y}\|_{2}^{2}+\frac{1}{3}\|\bm{H}_{0}\|_{2}\|\bm{y}\|_{2}^{3}$ (28b)
$\leq\|\bm{\tilde{d}}\|_{2}\|\bm{y}\|_{2}+\lambda_{1}\|\bm{y}\|_{2}^{2}+\frac{r^{\frac{3}{2}}\epsilon_{Q}}{3}\|\bm{y}\|_{2}^{3}.$ (28c)
In the second step, Eq. (28b), we replaced each term on the right-hand side with an upper bound on that term, and used the definition of a matrix 2-norm. In the last step, Eq. (28c), we used the Cauchy-Schwartz inequality and the matrix norm inequality $\|\bm{H}_{0}\|_{2}\leq\sqrt{r\times r^{2}}\|\bm{H}_{0}\|_{\text{max}}$, where $\|\cdot\|_{\text{max}}$ denotes the element-wise max of the matrix. For our upper bound estimate on $\dot{v}(\bm{y})$ to be negative, we consider solving $\|\bm{\tilde{d}}\|_{2}\|\bm{y}\|_{2}+\lambda_{1}\|\bm{y}\|_{2}^{2}+\frac{r^{\frac{3}{2}}\epsilon_{Q}}{3}\|\bm{y}\|_{2}^{3}=0$, from which it follows that
$\rho_{-}<\|\bm{y}\|_{2}<\rho_{+},$
$\rho_{+}$ $=\frac{3}{2r^{\frac{3}{2}}\epsilon_{Q}}\left(\sqrt{\lambda^{2}_{1}-\frac{4r^{\frac{3}{2}}\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}}-\lambda_{1}\right),$
$\rho_{-}$ $=-\frac{3}{2r^{\frac{3}{2}}\epsilon_{Q}}\left(\sqrt{\lambda^{2}_{1}-\frac{4r^{\frac{3}{2}}\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}}+\lambda_{1}\right).$
These equations can be manipulated into Equations (25) and (26). To guarantee $\rho_{+}$ and $\rho_{-}$ both being real, $\Delta=\lambda^{2}_{1}-\frac{4r^{\frac{3}{2}}\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}$ must be non-negative. For all $\bm{y}\in(\rho_{-},\rho_{+})$, we now have $v(\bm{y})>0$ and $\dot{v}(\bm{y})<0$, which yields the claimed result. ∎
@@ II.5 Interpretation of the local stability radius
Figure 2: Sketch of the trapping region. $\dot{v}<0$ is observed in $\bm{\Omega}_{\rho_{+}}\setminus\bm{\Omega}_{\rho_{-}}$. $\partial\bm{\Omega}_{\rho_{+}}$ has a complex geometry but can be bounded below by $\partial\bm{B}_{\rho_{+}}$. $\bm{\Omega}_{\rho_{-}}$ has $\dot{v}>0$, a shape close to ellipsoid, and can be bounded above by $\partial\bm{B}_{\rho_{-}}$. Therefore long-term stability is guaranteed for any initial conditions $\|\bm{y}(t_{0})\|\leq\rho_{+}$.
The $\rho_{+}$ and $\rho_{-}$ are two nontrivial roots of Eq. (28c). As shown in the Fig. 2, $\rho_{+}$ is a conservative estimate of the diameter of the inscribed ball $\textbf{B}_{\rho_{+}}$ in the locally stable region $\bm{\Omega}_{\rho_{+}}$. Meanwhile $\rho_{-}$ is an estimate of the diameter of the set $\Omega_{\rho_{-}}$ and $\dot{v}=0$ at the boundary $\partial\Omega_{\rho_{-}}$. Negative energy growth in $\textbf{B}_{\rho_{+}}\setminus\textbf{B}_{\rho_{-}}$ is guaranteed according to Thm. 2 and inside $\textbf{B}_{\rho_{-}}$ there exists a locally attracting trapping region $\bm{\Omega}_{\rho_{-}}$, inside of which positive energy growth can be observed. Then the long-time behavior of solutions to the system (22) with initial conditions $\bm{y}_{0}\in\textbf{B}_{\rho_{+}}$ alternates between $\textbf{B}_{\rho_{-}}\setminus\bm{\Omega}_{\rho_{-}}$ and $\bm{\Omega}_{\rho_{-}}$, as shown in the Fig.. 3. In a special case where $\bm{y}_{0}$ is a real solution to Eq. (28a), the long-time behavior of the system stays at the boundary of $\bm{\Omega}_{\rho_{-}}$. It is not difficult to conclude that
$\lim_{\epsilon_{Q}\to 0}\rho_{+},$ $=\infty\quad\lim_{\epsilon_{Q}\to 0}\rho_{-}=\frac{\|\tilde{\bm{d}}\|_{2}}{|\lambda_{1}(\tilde{\bm{A}}^{S})|},$ (29)
@@
$\dot{v}(\bm{y})=\bm{\tilde{d}}^{T}\bm{y}+\bm{y}^{T}\bm{\tilde{A}}^{S}\bm{y}$ $\leq\|\bm{\tilde{d}}\|_{2}\|\bm{y}\|_{2}+\lambda_{1}\|\bm{y}\|_{2}^{2}$
$=\|\bm{y}\|_{2}\left(\|\bm{\tilde{d}}\|_{2}+\lambda_{1}\|\bm{y}\|_{2}\right),$
from which we can see that for all $\|\bm{y}\|_{2}>\|\tilde{\bm{d}}\|_{2}/|\lambda_{1}|$, $\dot{v}<0$. And an estimate of the radius of the ball-shape trapping region is given by $R_{\bm{m}}=\|\tilde{\bm{d}}\|_{2}/|\lambda_{1}|$.
@@
$\epsilon_{Q}=\frac{3\lambda^{2}_{1}(\tilde{\bm{A}}^{S})}{4r^{\frac{3}{2}}\|\tilde{\bm{d}}\|_{2}},$ (30)
which provides a condition on the smallness of the totally symmetric part of the quadratic nonlinearity in order for this shell of stability to exist. We use this condition in Appendix A to evaluate the sensitivity of this stability estimate to small perturbations in the model coefficients and illustrate evidence that the problem of retaining stability in machine-learned quadratically nonlinear models is generally ill-conditioned.
Note that this condition becomes more stringent in higher dimensional state spaces, although there are also more parameters in $\bm{m}$ that may be able to make $\|\bm{d}\|_{2}$ smaller (the region near the origin gets more stable) or $\lambda^{2}_{1}(\bm{A}^{S})$ bigger (the unstable region where the non-energy-preserving quadratic nonlinearities begin to dominate gets shifted outwards). Note when $\|\bm{d}\|_{2}=0$, $\bm{\Omega}_{\rho_{-}}$ collapses to an asymptotically stable fixed point $\bm{m}$. Regardless, the scaling with $r^{\frac{3}{2}}$ merely reflects that we bounded $\|\bm{H}_{0}\|_{2}$ above by its max-norm but an equivalent Theorem 2 exists if we assume instead that, e.g. $\|\bm{H}_{0}\|_{F}$ is small, and then the scaling with $r$ vanishes.
@@ III Extended trapping SINDy algorithm
Table 1: Description of extended trapping SINDy hyperparameters.
$\gamma$
Specifies the strength of sparsity-promotion through the $l_{1}$ regularizer. For simplicity, all results in this paper use $\gamma=0$.
$\eta$
Specifies how strongly to push the algorithm towards models with Hurwitz $\tilde{\bm{A}}^{S}$.
$\beta$
Determines how large the deviation is from $\bm{Q}$ having zero totally symmetric part.
@@ III.1 Standard and trapping SINDy algorithm
where $\dot{\bm{X}}_{:,j}$ denotes the $j^{th}$ column of $\dot{\bm{X}}$, $\|\cdot\|_{0}$ counts the number of nonzero elements and $\gamma$ is related to a threshold value such that a new nonzero element is allowed to be added if it reduces the $\ell_{2}$ norm of the error by at least $\gamma$, which promotes the sparsity of $\bm{\Xi}$. One may also notice that $\|\cdot\|_{0}$ is not convex, which leads to a nonconvex optimization. One way to solve this is instead using $\ell_{1}$ norm resulting in the LASSO [31], which can be sparsity-promoting as well though it may be computationally expensive [32]. A computationally efficient alternative demonstrated by Brunton et al[23] is to implement the sequential threshold least-squares algorithm which allows a single parameter $\gamma$ to realize the desired degree of sparsity.
@@
$\bm{\xi}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{\phi}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (33)
$+\bm{\phi}^{T}(\bm{C}\bm{\xi}^{\prime}-\bm{q})\biggr{]}$
where $\bm{\hat{\Theta}}(\bm{X})=\bm{I}\otimes\bm{\Theta}(\bm{X})$, $\dot{\bm{X}}[:]$ is the vectorized $\dot{\bm{X}}$ so that we have $\bm{\hat{\Theta}}\in\mathbb{R}^{rM\times rN}$, $\dot{\bm{X}}[:]\in\mathbb{R}^{rM}$. The number of constraints of the energy-preserving structure is given by $p=r(r+1)(r+2)/6$ and therefore we have $\bm{C}\in\mathbb{R}^{p\times rN}$, $\bm{q}\in\mathbb{R}^{p}$ and the Lagrange multiplier $\bm{\phi}\in\mathbb{R}^{p}$. Additionally, for the linear equality constraints of the energy-preserving structure, $\bm{q}=0$ and the number of free parameters is given by $rN-p=2p$ [13]. If needed, these constraints can be used to impose any additional linear relationships from physical knowledge, such as enforcing global invariant preservation or enforcing a priori known values of any entries $\xi_{i}$.
It is still not enough to guarantee the global stability of the linear-quadratic systems with the energy-preserving constraint. Based on the Schlegel and Noack trapping theorem, Kaptanoglu et al[13] enforce global stability for quadratically nonlinear dynamics by searching for an $\bm{m}$ that makes $\tilde{\bm{A}}^{S}$ Hurwitz without breaking the energy-preserving structure of the systems. In other words, the largest real part of the eigenvalue of $\tilde{\bm{A}}^{S}$ should be negative, so that the optimization problem becomes,
$\bm{\xi}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{\phi},\bm{m}}$ $\Biggr{[}\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (34)
$\left.+\bm{\phi}^{T}(\bm{C}\bm{\xi}^{\prime}-\bm{q})+\frac{\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})}{\eta}\right],$
where the new term denotes that the largest eigenvalue of $\tilde{\bm{A}}^{S}$ will be enforced to be negative so that $\tilde{\bm{A}}^{S}$ is Hurwitz, and the magnitude of whose loss is modulated by the hyperparameter $\eta$. The details of the algorithm used for solving the optimization problem in (34) will not be introduced in this work; see Kaptanoglu et al[13] for more details.

### v3

@@ II.4 Analytical estimates of the local stability of linear-quadratic systems
The time derivative of the Lyapunov function given by Eq. (II.4) is
$\dot{v}(\bm{y})$ $=\bm{y}^{T}\bm{P}\dot{\bm{y}}$
$=\bm{d}^{T}\bm{P}\bm{y}+\frac{1}{2}\bm{y}^{T}(\bm{P}\bm{A}+\bm{A}^{T}\bm{P})\bm{y}+\bm{y}^{T}\bm{P}\bm{H}(y\otimes y)$
We can then introduce $\tilde{Q}^{\prime}_{ijk}=\tilde{Q}_{ijk}+\tilde{Q}_{jki}+\tilde{Q}_{kij}$, and $\bm{H}_{0}=\mathcal{T}(\tilde{\bm{Q}}^{\prime})$. See that
$\bm{y}\bm{H}_{0}(\bm{y}\otimes\bm{y})$ $=\sum_{i,j,k}\tilde{Q}^{\prime}_{ijk}y_{i}y_{j}y_{k}$
$=\sum_{i,j,k}(\tilde{Q}_{ijk}+\tilde{Q}_{jki}+\tilde{Q}_{kij})y_{i}y_{j}y_{k}$
$=3\sum_{i,j,k}\tilde{Q}_{ijk}y_{i}y_{j}y_{k}$
$=3\bm{y}^{T}\bm{P}\bm{H}(y\otimes y)$
Then by Eq. (26), $\dot{v}(\bm{y})$ can also be written as
$\dot{v}(\bm{z})$ $=\bm{\tilde{d}}^{T}\bm{z}+\bm{z}\bm{\tilde{A}}^{S}\bm{z}+\frac{1}{3}\bm{z}^{T}\bm{P}^{-\frac{1}{2}}\bm{H}_{0}(\bm{P}^{-\frac{1}{2}}\otimes\bm{P}^{-\frac{1}{2}})(\bm{z}\otimes\bm{z})$
$=\bm{\tilde{d}}^{T}\bm{z}+\bm{z}\bm{\tilde{A}}^{S}\bm{z}+\frac{1}{3}\bm{z}^{T}\bm{\tilde{H}}_{0}(\bm{z}\otimes\bm{z})$
Here $\bm{\tilde{d}}=\bm{P}^{\frac{1}{2}}\bm{d}$, $\bm{\tilde{A}}^{S}=\frac{1}{2}\left(\bm{P}^{\frac{1}{2}}\bm{A}\bm{P}^{-\frac{1}{2}}+\bm{P}^{-\frac{1}{2}}\bm{A}^{T}\bm{P}^{\frac{1}{2}}\right)$ and $\bm{\tilde{H}}_{0}=\bm{P}^{-\frac{1}{2}}\bm{H}_{0}(\bm{P}^{-\frac{1}{2}}\otimes\bm{P}^{-\frac{1}{2}})$. We next consider the region where its derivative is negative. It follows that
$\dot{v}(\bm{z})$ $=\bm{\tilde{d}}^{T}\bm{z}+\bm{z}\bm{\tilde{A}}^{S}\bm{z}+\frac{1}{3}\bm{z}^{T}\bm{\tilde{H}}_{0}(\bm{z}\otimes\bm{z})$ (31a)
$\leq\|\bm{\tilde{d}}^{T}\bm{z}\|_{2}+\lambda_{1}\|\bm{z}\|_{2}^{2}+\frac{1}{3}\|\bm{\tilde{H}}_{0}\|_{2}\|\bm{z}\|_{2}^{3}$ (31b)
$\leq\|\bm{\tilde{d}}\|_{2}\|\bm{z}\|_{2}+\lambda_{1}\|\bm{z}\|_{2}^{2}+\frac{\epsilon_{Q}}{3}\|\bm{z}\|_{2}^{3}.$ (31c)
In the second step, Eq. (31b), we replaced each term on the right-hand side with an upper bound on that term, and used the definition of a matrix 2-norm. In the last step, Eq. (31c), we used the Cauchy-Schwartz inequality and the matrix norm inequality $\|\bm{\tilde{H}}_{0}\|_{2}\leq\|\bm{\tilde{H}}_{0}\|_{F}$, where $\|\cdot\|_{F}$ denotes the Frobenius norm of the matrix. For our upper bound estimate on $\dot{v}(\bm{z})$ to be negative, we consider solving $\|\bm{\tilde{d}}\|_{2}\|\bm{z}\|_{2}+\lambda_{1}\|\bm{z}\|_{2}^{2}+\frac{\epsilon_{Q}}{3}\|\bm{z}\|_{2}^{3}=0$, from which it follows that
$\rho_{-}<\|\bm{z}\|_{2}<\rho_{+},$
$\rho_{+}$ $=\frac{3}{2\epsilon_{Q}}\left(\sqrt{\lambda^{2}_{1}-\frac{4\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}}-\lambda_{1}\right),$
$\rho_{-}$ $=-\frac{3}{2\epsilon_{Q}}\left(\sqrt{\lambda^{2}_{1}-\frac{4\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}}+\lambda_{1}\right).$
These equations can be manipulated into Equations (28) and (29). To guarantee $\rho_{+}$ and $\rho_{-}$ both being real, $\Delta=\lambda^{2}_{1}-\frac{4\epsilon_{Q}}{3}\|\bm{\tilde{d}}\|_{2}$ must be non-negative. For all $\|\bm{y}\|_{\bm{P}}=\|\bm{z}\|_{2}\in(\rho_{-},\rho_{+})$, we now have $v(\bm{y})>0$ and $\dot{v}(\bm{y})<0$, which yields the claimed result. ∎
@@ II.5 Interpretation of the local stability radius
The $\rho_{+}$ and $\rho_{-}$ are two nontrivial roots of Eq. (31c). As shown in the Fig. 2, $\rho_{+}$ is a conservative estimate of the diameter of the inscribed ball $\textbf{B}_{\rho_{+}}$ in the locally stable region $\bm{\Omega}_{\rho_{+}}$. Meanwhile $\rho_{-}$ is an estimate of the diameter of the set $\Omega_{\rho_{-}}$ and $\dot{v}=0$ at the boundary $\partial\Omega_{\rho_{-}}$. Negative energy growth in $\textbf{B}_{\rho_{+}}\setminus\textbf{B}_{\rho_{-}}$ which is an elliptical annulus in the state space of $\bm{y}$, is guaranteed according to Thm. 2 and inside $\textbf{B}_{\rho_{-}}$ there exists a region $\bm{\Omega}_{\rho_{-}}$, inside of which positive energy growth can be observed. Then the long-time behavior of solutions to the system (23) with initial conditions $\bm{P}^{\frac{1}{2}}\bm{y}_{0}=\bm{z}_{0}\in\textbf{B}_{\rho_{+}}$ can alternate between $\textbf{B}_{\rho_{-}}\setminus\bm{\Omega}_{\rho_{-}}$ and $\bm{\Omega}_{\rho_{-}}$, as shown in the Fig. 3. In a special case where $\bm{z}_{0}$ is a real solution to Eq. (31a), the long-time behavior of the system stays at the boundary of $\bm{\Omega}_{\rho_{-}}$. It is not difficult to conclude that
$\lim_{\epsilon_{Q}\to 0}\rho_{+}$ $=\infty,\quad\lim_{\epsilon_{Q}\to 0}\rho_{-}=\frac{\|\tilde{\bm{d}}\|_{2}}{|\lambda_{1}(\tilde{\bm{A}}^{S})|},$ (32)
@@
$\dot{v}(\bm{z})=\bm{\tilde{d}}^{T}\bm{z}+\bm{z}^{T}\bm{\tilde{A}}^{S}\bm{z}$ $\leq\|\bm{\tilde{d}}\|_{2}\|\bm{z}\|_{2}+\lambda_{1}\|\bm{z}\|_{2}^{2}$
$=\|\bm{z}\|_{2}\left(\|\bm{\tilde{d}}\|_{2}+\lambda_{1}\|\bm{z}\|_{2}\right),$
from which we can see that for all $\|\bm{z}\|_{2}>\|\tilde{\bm{d}}\|_{2}/|\lambda_{1}|$, $\dot{v}<0$. And an estimate of the radius of the ball-shape trapping region is given by $R_{\bm{m}}=\|\tilde{\bm{d}}\|_{2}/|\lambda_{1}|$.
Figure 2: Sketch of the trapping region. $\dot{v}<0$ is observed in $\bm{\Omega}_{\rho_{+}}\setminus\bm{\Omega}_{\rho_{-}}$. $\partial\bm{\Omega}_{\rho_{+}}$ has a complex geometry but can be bounded below by $\partial\bm{B}_{\rho_{+}}$. $\bm{\Omega}_{\rho_{-}}$ has $\dot{v}>0$, a shape close to ellipsoid, and can be bounded above by $\partial\bm{B}_{\rho_{-}}$. Therefore long-term stability is guaranteed for any initial conditions $\|\bm{y}(t_{0})\|_{\bm{P}}=\|\bm{z}(t_{0})\|_{2}\leq\rho_{+}$.
@@
$\epsilon_{Q}=\frac{3\lambda^{2}_{1}(\tilde{\bm{A}}^{S})}{4\|\tilde{\bm{d}}\|_{2}},$ (33)
which provides a condition on the smallness of the totally symmetric part of the quadratic nonlinearity in order for this shell of stability to exist. We use this condition in Appendix A to evaluate the sensitivity of this stability estimate to small perturbations in the model coefficients and illustrate evidence that the problem of retaining stability in machine-learned quadratically nonlinear models is generally ill-conditioned. Note when $\|\bm{d}\|_{2}=0$, $\bm{\Omega}_{\rho_{-}}$ collapses to an asymptotically stable fixed point $\bm{m}$.
@@ III.1 Standard and trapping SINDy algorithm
where $\dot{\bm{X}}_{:,j}$ denotes the $j^{th}$ column of $\dot{\bm{X}}$, $\|\cdot\|_{0}$ counts the number of nonzero elements and $\gamma$ determines the strength of sparsity-promotion during identification of the $\bm{\Xi}$. One may also notice that $\|\cdot\|_{0}$ is not convex, which leads to a nonconvex optimization. One way to ameliorate this is to instead use the $\ell_{1}$ norm,resulting in the LASSO [39]. The $\ell_{1}$ is convex but promotes a weaker form of sparsity. A computationally efficient alternative to directly solving the $\ell_{0}$-regularized problem was demonstrated in Brunton et al[28], where a sequential threshold least-squares algorithm was used to greedily solve the problem; its convergence properties were proven in [40].
@@
$\bm{\xi}=\operatorname*{argmin}_{\bm{\xi}^{\prime}}\max_{\bm{\zeta}}\biggr{[}$ $\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (36)
$+\bm{\zeta}^{T}(\bm{C}\bm{\xi}^{\prime}-\bm{q})\biggr{]}$
where $\bm{\hat{\Theta}}(\bm{X})=\bm{I}\otimes\bm{\Theta}(\bm{X})$, $\dot{\bm{X}}[:]$ is the vectorized $\dot{\bm{X}}$ so that we have $\bm{\hat{\Theta}}\in\mathbb{R}^{rM\times rN}$, $\dot{\bm{X}}[:]\in\mathbb{R}^{rM}$. The number of constraints of the energy-preserving structure is given by $p=r(r+1)(r+2)/6$ and therefore we have $\bm{C}\in\mathbb{R}^{p\times rN}$, $\bm{q}\in\mathbb{R}^{p}$ and the Lagrange multiplier $\bm{\zeta}\in\mathbb{R}^{p}$. Additionally, for the linear equality constraints of the energy-preserving structure, $\bm{q}=0$ and the number of free parameters is given by $rN-p=2p$. If needed, these constraints can be used to impose any additional linear relationships from physical knowledge, such as enforcing global invariant preservation or enforcing a priori known values of any entries $\xi_{i}$.
It is still not enough to guarantee the global stability of the linear-quadratic systems with the energy-preserving constraint. Based on the Schlegel and Noack trapping theorem, enforcing global stability for quadratically nonlinear dynamics requires searching for an $\bm{m}$ that makes $\tilde{\bm{A}}^{S}$ Hurwitz without breaking the energy-preserving structure of the systems. In other words, the largest real part of the eigenvalue of $\tilde{\bm{A}}^{S}$ should be negative, so that the optimization problem becomes,
$\bm{\xi},\bm{\zeta},\bm{m}=\operatorname*{argmin}_{\bm{\xi}^{\prime},\bm{m}}$ $\max_{\bm{\zeta}}\Biggr{[}\frac{1}{2}\|\dot{\bm{X}}[:]-\bm{\hat{\Theta}}(\bm{X})\bm{\xi}^{\prime}\|^{2}_{2}+\gamma\|\bm{\xi}^{\prime}\|_{1}$ (37)
$\left.+\bm{\zeta}^{T}(\bm{C}\bm{\xi}^{\prime}-\bm{q})+\frac{\bm{\lambda}_{1}(\tilde{\bm{A}}^{S})}{\eta}\right],$
where the new term denotes that the largest eigenvalue of $\tilde{\bm{A}}^{S}$ will be enforced to be negative so that $\tilde{\bm{A}}^{S}$ is Hurwitz, and the magnitude of whose loss is modulated by the hyperparameter $\eta$. One should notice that the last term of this problem is not convex but convex composite, making the problem non-convex, however the details of the algorithm used for solving the optimization problem in (37) will not be introduced in this work; see Kaptanoglu et al[2] for more details.